OpenStudy (anonymous):

Simplify: (–5) • (7) – (6) • (–2) A. –23 B. –35 C. 10 D. 47

OpenStudy (anonymous):

-5 * 7 = -35 -6 * -2 = 12 -35 + 12 = -23

OpenStudy (anonymous):

Good job @jadeybells

OpenStudy (anonymous):

Um.. thanks?

OpenStudy (anonymous):

Thank you! :D